OpenStudy (anonymous):

if f(x) is a monic quatric polynomial such that f(-1)=-1, f(2)=-4, f(-3)=-9, andf(4)=-16 then f(1)=

OpenStudy (anonymous):

you need to create an arbitrary monic quartic polynomial: \[f(x) = x^4+bx^3+cx^2+dx+e\] plug in the values, and solve the resulting system of equations for b, c, d, and e.

OpenStudy (anonymous):

you end up with:\[f(x) = x^4-2x^3-14x^2+14x+24\]

OpenStudy (anonymous):

oh, so to finish the equation, f(1) would be: \[f(1) = 1-2-14+14+24 = 23\]

OpenStudy (anonymous):

thank you